OpenStudy (anonymous):

Write an expression that can be used to solve the problem. During practice, it took Shawn 27 min to swim 81 laps. If Shawn swam at a constant rate, how long did it take him to complete each lap? A.27*81 B.27+81 C.81/27 D.27/81

OpenStudy (anonymous):

81 / 27 is to find the answer, so it's C.

OpenStudy (anonymous):

oh ok ty can yuh help me with 4 more or no?

OpenStudy (anonymous):

Yeah, sure.

OpenStudy (anonymous):

ok ty hold on

OpenStudy (anonymous):

which expression could be used to solve the problem? While preparing a recipe, Chris noticed that 4 cans held a total of 34 oz of tomato sauce. The cans were all the same size. How many ounces did each can hold? A.34*4 B.34+4 C.34/4 D.4/34

OpenStudy (anonymous):

@ Dalycia @greenlegodude57 is very wrong. The answer is D. 27/81. We are looking for time and not distance.

OpenStudy (anonymous):

34 / 4

OpenStudy (anonymous):

@Perculiar Actually, no I'm right. To check your answer 3 MINUTES TIMES THE TOTAL 27 MINUTES should get you back to 81. 3 x 27 = 81.

OpenStudy (anonymous):

I got 3 minutes from 81 / 27.

OpenStudy (anonymous):

@Dalycia Keep posting your questions..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold on sorry

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

168 / 8.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@greenlegodude57 @Dalycia the question is it took 25mins to swam 81 lap. how many minutes will it take to for each lap. i.e 27mins=81laps therefore: ans.D = 27/81x1lap = 0.33secs on every lap. Check answer (0.333secs x 81laps) = 26.973mins roundup =27mins

OpenStudy (anonymous):

So its not D.?

OpenStudy (anonymous):

First find out what is the cost of one notebook: 15 / 5 = 3. Since it's $3.00 for one notebook, multiply 3 with 8 to find out the value of 8 notebooks: 8 x 3 = $24.00 for 8 notebooks.

OpenStudy (anonymous):

No, the answer is NOT D. It's C.

OpenStudy (anonymous):

The answer to the notebook one is D, though.

OpenStudy (anonymous):

The answer is D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I have already explained it to you, Perculiar.

OpenStudy (anonymous):

After @Dalycia submits her test, we'll see who's right.

OpenStudy (anonymous):

D. $24

OpenStudy (anonymous):

Ok oooo @greenlegodude57 @Dalycia

OpenStudy (anonymous):

15 / 10.

OpenStudy (anonymous):

I missed 1

OpenStudy (anonymous):

Which one..?

OpenStudy (anonymous):

the last 1

OpenStudy (anonymous):

Ohh, sorry I thought it said bottles per gallon instead of gallons per bottle. So the answer was 10 / 15.

OpenStudy (anonymous):

its ok ty for ur help btw

OpenStudy (anonymous):

Np, I was right on that swimming one, @Perculiar

OpenStudy (anonymous):

@greenlegodude57 plz check the first question again. how come u r getting other questions correctly but still fails it - the swimming

OpenStudy (anonymous):

@Perculiar She said she got that one right, but she got the last one wrong.